# Learning Reward Machines from Partially Observed Policies

### v2

@@ 1 Introduction
As an extension to inverse reinforcement learning (IRL) (Ng et al., 2000), one could ask the question of learning RMs, which enables agents to autonomously extract structured representations of tasks, significantly enhancing their ability to solve complex, temporally extended problems. By learning reward machines directly from demonstrations, agents can adapt to tasks without requiring manually specified task representations, making this approach scalable and practical for real-world applications, such as robotic manipulation and autonomous vehicle navigation (Camacho et al., 2019; Icarte et al., 2023; Xu et al., 2020; Baert et al., 2024; Camacho et al., 2021). This capability is especially valuable in environments where high-level task-relevant features (aka, atomic propositions) are observable, underscoring the importance of learning RMs in advancing autonomous decision-making systems. Beyond autonomy applications, IRL and RM learning can also be used to infer the agent’s (e.g., humans’) intentions to design incentives or better decision making environments (Nitschke et al., 2024).
@@ 2.1 Markov Decision Processes and Reward Machines
$\mathrm{Reach}(\sigma)$ =$=\{s\in\mathcal{S}\mid\tau\in\mathcal{T}_{s}^{\mathrm{fin}}\text{s.t. }L(\tau)=\sigma,\tau_{|\tau|}=s\}.$
A Reward Machine (RM) is a tuple $\mathcal{R}=(\mathcal{U},u_{I},\mathrm{AP},\delta_{\mathbf{u}},\delta_{\mathbf{r}})$ which consists of a finite set of states $\mathcal{U}$, an initial state $u_{I}\in\mathcal{U}$, an input alphabet $\mathrm{AP}$, a (deterministic) transition function $\delta_{\mathbf{u}}:\mathcal{U}\times\mathrm{AP}\to\mathcal{U}$, and an output function $\delta_{\mathbf{r}}:\mathcal{U}\times\mathrm{AP}\to\mathbb{R}$. To avoid ambiguity between MDP states and RM states, the later will be referred to as nodes. The reward machine without the reward is denoted as $\mathcal{G}\triangleq\mathcal{R}/\delta_{\mathbf{r}}$, and we refer to it as a reward machine model. We extend the definition of the transition function to define $\delta_{\mathbf{u}}^{*}:\mathcal{U}\times(\mathrm{AP})^{*}\to\mathcal{U}$ as $\delta_{\mathbf{u}}^{*}(u,l_{0},\cdots,l_{k})=\delta_{\mathbf{u}}(\cdots(\delta_{\mathbf{u}}(\delta_{\mathbf{u}}(u,l_{0}),l_{1}),\cdots,l_{k})$. Given a state $u\in\mathcal{U}$, we define the paths of $u$ as the input words which can be used to reach $u$:
@@ 3.1 SAT Encoding
We encode the RM learning problem into a SAT problem. Specifically, we use SAT to encode a graph with $n\leq u_{\mathrm{max}}$ nodes and associate a Boolean variable with each edge in the graph. Each node has $|\mathrm{AP}|$ outgoing edges. We define the Boolean variables $\{b_{ikj}\mid 1\leq i,j\leq n,1\leq k\leq|\mathrm{AP}|\}$ as:
[…]
3.1.1 Determinism
@@ 3.1.2 Full Specification
We can combine the conditions of Sections 3.1.1 and 3.1.2 into one condition on each $B_{k}$ enforcing that each row has exactly one entry with value $1$.
3.1.3 Negative Examples
@@ 3.2 Algorithm
We start with one node and increases the number $n$ of nodes until the following SAT problem is feasible when instantiated for a graph with $n$ nodes:
@@ 4       $n\leftarrow n+1$
9$(\mathcal{M}/r)^{\mathrm{learned}}\leftarrow\mathcal{G}^{\mathrm{learned}}\times\mathcal{M}_{L}$
@@
It is important to emphasize that up to line 7 in Algorithm 1, neither the optimality of the prefix tree policy nor the transition kernel $\mathcal{P}$ of the MDP is used. That is, our method learns a reward machine model $\mathcal{G}^{\mathrm{learned}}$ in a model-free fashion as long as the prefix-tree policy is induced by a reward machine. The optimality with respect to the MaxEnt objective in (1) and the transition kernel only comes into play to extract the numerical reward values in lines 8-11.
@@ 4.2 MAX-SAT: SAT with Robustness to Incorrect Negative Examples
Even when the set of negative examples is constructed from pairs of prefixes for which we have a high confidence of being true negative examples, some pairs may still be mislabeled as negative examples. We refer to those as false positives. To deal with these false positive, we implement a weighted MAX-SAT (Manquinho et al., 2009; Biere et al., 2009) variant of the SAT problem. Concretely, the weighted MAX-SAT problem consists of finding a Boolean assignment of the variables $b_{ikj}$ such that (i) constraints (5), (6) and (9) hold, i.e., the recovered RM is deterministic, non-stuttering, and fully specified; and (ii) the constraint (8) holds for a maximum number of pairs $\{\sigma,\sigma^{\prime}\}$ in $\mathcal{E}^{-}_{l}$. In other words, (5), (6) and (9) are hard constraints whereas the constraints coming from (8) are soft constraints. This approach allows for robustness to misclassifications of negative examples. We empirically demonstrate in Section 5.3 that this weighted MAX-SAT formulation consistently infers the smallest set of consistent reward machine models, even in the presence of false positives.
@@ 5 Experiments
To demonstrate the generality and efficiency of our approach, we apply it to a diverse set of domains, from classical grid-based MDPs to a continuous robotic control task and a real-world biological navigation dataset. In every experiment, we fix the discount factor to $\gamma=0.99$ and the regularization weight to $\lambda=1.0$ when solving Problem (2), both for generating demonstration traces and for reward recovery. Our code is implemented in Python, and the Z3 library (De Moura and Bjørner, 2008) is used for solving the SAT and weighted MAX-SAT problems. To enumerate all the satisfying assignments of the SAT problem, we add a constraint that the next solution should be different every time our SAT solver finds a solution. We will release the complete implementation on GitHub, upon acceptance to adhere with the double-blind review process.
@@ 5.1 Tabular GridWorld MDPs
Our first experiment (patrol) is on the running-example patrol task of Figure 1. By setting $u_{\mathrm{max}}=4$, our bound from Theorem 1 is $256$. However, with only a depth-$6$ prefix tree policy, we end up with a total of $\mathbf{6}$ solutions. These are all the possible renamings of the true reward machine (see Figure 1(b)), meaning that the true reward machine is learned up-to-renaming with a depth-$6$ prefix tree policy. Table 1 summarizes the findings, where we also show how the non-stuttering condition of Section 3.1.4 helps reduce the size of the negative example set, yet still recovering the true reward machine model. While some reduction in the SAT solver time is achieved, the drastic gain is in the time required to encode all the negative examples into the SAT solver, making the overall procedure orders of magnitude faster. For the remaining experiments, non-stuttering is assumed.
@@
patrol with (9) 6 16 4 7264 0.81
patrol without (9) 6 16 4 168341 6.97
@@
For our second experiment (patrol-hallway), we add a hallway between the rooms, as shown in Figure 1(c). This is a $3\times 3$ grid world, where the corresponding label of each room is shown. The reward machine is kept the same, and the atomic proposition $\mathrm{H}$ is added as a self-loop to all the nodes. With this added hallway, longer atomic proposition prefixes are required to reach all nodes of the reward machine, showing how the underlying MDP affects the required depth for learning a reward machine. For example, the shortest atomic proposition that can reach $u_{3}$ is now $\sigma=\mathrm{AHBHC}$ instead of $\mathrm{ABC}$ as in the previous example. With a depth-$9$ prefix tree policy, the reward machine is again learned up-to-renaming.
@@ 5.2 Tabular BlockWorld MDPs
The setup for this experiment is a modified block world problem (Khodeir et al., 2023; Wolfe and Barto, 2006). There are three blocks colored green, yellow and red, as well as 3 piles. Each stacking configuration of the blocks is a state of the MDP, and the action space consists of selecting a pile to pick from and a pile to place onto. We can only grab the top block of any stack. Action outcomes are assumed deterministic. The goal in the first task (stack) is to stack the blocks in the ordered stacking configurations $\mathbf{st_{1}},\mathbf{st_{2}},\mathbf{st_{3}}$, shown in Figure 2(a). All other states have the label $\mathbf{i}$, denoting intermediate states. The corresponding reward machine is shown in Figure 2(b). If the robot stacks the blocks in the order $\mathbf{st_{1}}\to\mathbf{st_{2}}\to\mathbf{st_{3}}$, it gets a reward of $1.0$. With a depth-$9$ prefix tree policy and $u_{\mathrm{max}}=3$, our algorithm recovers $\mathbf{2}$ consistent reward machines, which are the true reward machine up-to-renaming. The findings are summarized in Table 2. Results obtained from a finite set of demonstrations are reported in Appendix B.1.1.
@@
For our second task (stack-avoid), we introduce a “bad” state, shown in Figure 3(a). The true reward machine is shown in Figure 3(b). The robot’s task is to stack the blocks in the order $\mathbf{st_{1}}\to\mathbf{st_{2}}$ without going through ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}}$. If it does so, it reaches $u_{2}$ and gets a reward of $1$ forever. If during execution it passes through ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}}$, it will get a smaller (yet more immediate) reward of $0.2$, but it will get stuck at $u_{3}$ with 00 reward forever. We note that the product policy is uniformly random in both $u_{2}$ and $u_{3}$. This means that proposition traces such as $\mathbf{st_{1},i,st_{2}}$ and $\mathbf{st_{1},i,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}},st_{2}}$ look identical from a policy perspective, as both reach nodes with uniformly random policies, while the first being more desirable than the second. By setting $u_{\mathrm{max}}=3$, a depth-$8$ policy was sufficient to recover the reward machine shown in Figure 3(c) up-to-renaming; that is, we find a smaller reward machine consistent for this task. Notably, setting $u_{\mathrm{max}}=4$ with the same depth policy yields more than a 1000 solutions, indicating that uniquely recovering the same sized reward machine -as the ground truth machine- requires longer depth policies. Results obtained from a finite set of demonstrations are reported in Appendix B.1.2.
@@
stack 9 60 9 31656 0.558
stack-avoid 8 60 9 24763 0.483
@@
5.3 2D-Continous Reacher Robotic Arm
Our setup for this experiment is a modified Reacher-v5 environment (Gymnasium (Towers et al., 2024); MuJoCo (Todorov et al., 2012)), where a 2D robotic arm must reach targets randomly placed within the arena. The agent’s state is a 10-D continuous vector consisting of the end-effector’s position and velocity, the target’s coordinates, and their mutual distance. Actions are continuous torques in $[-1,1]^{2}$ applied at the elbow and shoulder joints.
Inspired by (Araki et al., 2021), we fix three colored targets — blue ($\mathbf{b}$), red ($\mathbf{r}$), and yellow ($\mathbf{y}$) — in distinct quadrants of the arena (see Figure 4). The goal is to reach them in the order $\mathbf{b}\to\mathbf{y}\to\mathbf{r}$. To train an expert policy, we discretize each torque dimension into five levels $\{-1.0,-0.5,0.0,0.5,1.0\}$ (yielding 25 total actions). We then employ Proximal Policy Optimization (PPO) (Schulman et al., 2017), as implemented in Stable-Baselines3 (Raffin et al., 2021), to maximize a reward given by the negative Euclidean distance between the end-effector and the active target.
@@
Table 3 reports, for various probability thresholds $\alpha$ (defined in Section 4.2), the size of the negative-example set $\mathcal{E}^{-}$ and its False Positive Rate (FPR) [^1] FPR is simply the number of false positives divided by $|\mathcal{E}^{-}|$.. As expected from Section 4, solving our SAT problem with all the negative examples yields $\mathbf{0}$ solutions. Instead, we employ our weighted MAX-SAT variant and recover two viable solutions (Figures 4(b) and 4(c)) up-to-renaming, which represent the best achievable given our dataset. These are the same solutions that we recover if instead we supervised with the ground-truth reward machine to remove the false positives, emphasizing the robustness of this approach.

### v3

@@ 1 Introduction
As an extension to inverse reinforcement learning (IRL) (Ng and Russell, 2000), one could ask the question of learning RMs, which enables agents to autonomously extract structured representations of tasks, significantly enhancing their ability to solve complex, temporally extended problems. By learning reward machines directly from demonstrations, agents can adapt to tasks without requiring manually specified task representations, making this approach scalable and practical for real-world applications, such as robotic manipulation and autonomous vehicle navigation (Camacho et al., 2019; Icarte et al., 2023; Xu et al., 2020; Baert et al., 2024; Camacho et al., 2021). This capability is especially valuable in environments where high-level task-relevant features (aka, atomic propositions) are observable, underscoring the importance of learning RMs in advancing autonomous decision-making systems. For instance, in a high-level indoor navigation or patrolling task, semantic room labels can act as such propositions. Beyond autonomy applications, IRL and RM learning can also be used to infer the agent’s (e.g., humans’) intentions to design incentives or better decision making environments (Nitschke et al., 2024). As will be demonstrated, such intent inference can also be used in neuroscience to analyze animal behavior and decision making.
@@ 2.1 Markov Decision Processes and Reward Machines
$\mathrm{Reach}(\sigma)$ $=\{s\in\mathcal{S}\mid\tau\in\mathcal{T}_{s}^{\mathrm{fin}}\text{ s.t. }L(\tau)=\sigma,\tau_{|\tau|}=s\}.$
A Reward Machine (RM) is a tuple $\mathcal{R}=(\mathcal{U},u_{I},\mathrm{AP},\delta_{\mathbf{u}},\delta_{\mathbf{r}})$ which consists of a finite set of states $\mathcal{U}$, an initial state $u_{I}\in\mathcal{U}$, an input alphabet $\mathrm{AP}$, a (deterministic) transition function $\delta_{\mathbf{u}}:\mathcal{U}\times\mathrm{AP}\to\mathcal{U}$, and an output function $\delta_{\mathbf{r}}:\mathcal{U}\times\mathrm{AP}\to\mathbb{R}$. To avoid ambiguity between MDP states and RM states, the latter will be referred to as nodes. The reward machine without the reward is denoted as $\mathcal{G}\triangleq\mathcal{R}/\delta_{\mathbf{r}}$, and we refer to it as a reward machine model. We extend the definition of the transition function to define $\delta_{\mathbf{u}}^{*}:\mathcal{U}\times(\mathrm{AP})^{*}\to\mathcal{U}$ as $\delta_{\mathbf{u}}^{*}(u,l_{0},\cdots,l_{k})=\delta_{\mathbf{u}}(\cdots(\delta_{\mathbf{u}}(\delta_{\mathbf{u}}(u,l_{0}),l_{1}),\cdots,l_{k})$. Given a state $u\in\mathcal{U}$, we define the paths of $u$ as the input words which can be used to reach $u$:
@@ 3.1 SAT Encoding
We encode the RM learning problem into a Boolean Satisfiability problem (SAT). SAT is the problem of determining whether there exists an assignment to variables of a given Boolean formula that makes it evaluate to true. While SAT is known to be NP-complete (Cook, 1971), there are several powerful off-the-shelf solvers capable of solving large practical instances (Biere et al., 2009). Specifically, we use SAT to encode a graph with $n\leq u_{\mathrm{max}}$ nodes and associate a Boolean variable with each edge in the graph. Each node has $|\mathrm{AP}|$ outgoing edges. We define the Boolean variables $\{b_{ikj}\mid 1\leq i,j\leq n,1\leq k\leq|\mathrm{AP}|\}$ as:
$b_{ikj}=\begin{cases}1\quad\text{if }i\overset{k}{\to}j,\\ 0\quad\text{Otherwise,}\end{cases}$ (4)
where we use the shorthand $i\overset{k}{\to}j$ to denote that proposition $k$ transitions node $i$ to node $j$, i.e., $\delta_{\mathbf{u}}(u_{i},k)=u_{j}$. We can encode several properties of the RM into Boolean constraints. Without loss of generality, we set node $1$ of the graph to be $u_{I}$. To make the derivation easier, we define for each atomic proposition $k$ an adjacency matrix $B_{k}$ with $(B_{k})_{ij}=b_{ikj}$. The Boolean constraints we add are due to determinism, full-specification, negative examples and non-stuttering of the learned reward machine. We expand on each of them below.
Determinism:
@@
We can combine the conditions of Sections 3.1 and 3.1 into one condition on each $B_{k}$ enforcing that each row has exactly one entry with value $1$.
Negative Examples:
@@ 3.2 Algorithm
We start with one node and increase the number $n$ of nodes until the following SAT problem is feasible when instantiated for a graph with $n$ nodes:
@@ 4 $n\leftarrow n+1$
[…]
@@
It is important to emphasize that up to line 7 in Algorithm 1, neither the optimality of the prefix tree policy nor the transition kernel $\mathcal{P}$ of the MDP is used. That is, our method learns a reward machine model $\mathcal{G}^{\mathrm{learned}}$ in a model-free fashion as long as the prefix-tree policy is induced by a reward machine. This is further illustrated through an example in Appendix C. The optimality with respect to the MaxEnt objective in (1) and the transition kernel only comes into play to extract the numerical reward values in lines 8-11.
@@ 4.2 MAX-SAT: SAT with Robustness to Incorrect Negative Examples
Even when the set of negative examples is constructed from pairs of prefixes for which we have a high confidence of being true negative examples, some pairs may still be mislabeled as negative examples. We refer to those as false positives. To deal with these false positives, we implement a weighted MAX-SAT (Manquinho et al., 2009; Biere et al., 2009) variant of the SAT problem. Concretely, the weighted MAX-SAT problem consists of finding a Boolean assignment of the variables $b_{ikj}$ such that (i) constraints (5), (6) and (9) hold, i.e., the recovered RM is deterministic, non‑stuttering, and fully specified; and (ii) the constraint (8) holds for a maximum number of pairs $\{\sigma,\sigma^{\prime}\}$ in $\mathcal{E}^{-}_{l}$. In other words, (5), (6) and (9) are hard constraints whereas the constraints coming from (8) are soft constraints with equal unitary weights. This approach allows for robustness to misclassifications of negative examples. We empirically demonstrate in Section 5.3 that this weighted MAX-SAT formulation consistently infers the smallest set of consistent reward machine models, even in the presence of false positives.
@@ 5 Experiments
To demonstrate the generality and efficiency of our approach, we apply it to a diverse set of domains, from classical grid‐based MDPs to a continuous robotic control task and a real‐world biological navigation dataset. In every experiment, we fix the discount factor to $\gamma=0.99$ and the regularization weight to $\lambda=1.0$ when solving Problem (2), both for generating demonstration traces and for reward recovery. Our code is implemented in Python, and the Z3 library (De Moura and Bjørner, 2008) is used for solving the SAT and weighted MAX-SAT problems. To enumerate all the satisfying assignments of the SAT problem, we add a constraint that the next solution should be different every time our SAT solver finds a solution. Our implementation code is made publicly available here: https://github.com/mlshehab/learning_reward_machines.git.
@@ 5.1 Tabular GridWorld MDPs
Our first experiment (patrol) is on the running-example patrol task of Figure 1. By setting $u_{\mathrm{max}}=4$, our bound from Theorem 1 is $256$. However, with only a depth-$6$ prefix tree policy, we end up with a total of $\mathbf{6}$ solutions. These are all the possible renamings of the true reward machine (see Figure 1(b)), meaning that the true reward machine is learned up-to-renaming with a depth-$6$ prefix tree policy. Table 1 summarizes the findings, where we also show how the non-stuttering condition of Section 3.1 helps reduce the size of the negative example set, yet still recovering the true reward machine model. While some reduction in the SAT solver time is achieved, the drastic gain is in the time required to encode all the negative examples into the SAT solver, making the overall procedure orders of magnitude faster. For the remaining experiments, non-stuttering is assumed.
@@
patrol with (9) 6 16 4 3076 0.51
patrol without (9) 6 16 4 30573 1.73
@@
For our second experiment (patrol-hallway), we add a hallway between the rooms, as shown in Figure 1(c). This is a $3\times 3$ grid world, where the corresponding label of each room is shown. The reward machine is kept the same, and the atomic proposition $\mathrm{H}$ is added as a self-loop to all the nodes. With this added hallway, longer atomic proposition prefixes are required to reach all nodes of the reward machine, showing how the underlying MDP affects the required depth for learning a reward machine. For example, the shortest atomic proposition that can reach $u_{3}$ is now $\sigma=\mathrm{AHBHC}$ instead of $\mathrm{ABC}$ as in the previous example. With a depth-$9$ prefix tree policy, the reward machine is again learned up-to-renaming. Additional experiments studying transferability and the sensitivity to the upper bound $u_{\mathrm{max}}$ can be found in appendices˜D and E.
@@ 5.2 Tabular BlockWorld MDPs
The setup for this experiment is a modified block world problem (Khodeir et al., 2023; Wolfe and Barto, 2006). There are three blocks colored green, yellow and red, as well as 3 piles. Each stacking configuration of the blocks is a state of the MDP, and the action space consists of selecting a pile to pick from and a pile to place onto. We can only grab the top block of any stack. Action outcomes are assumed deterministic. The goal in the first task (stack) is to stack the blocks in the ordered stacking configurations $\mathbf{st_{1}},\mathbf{st_{2}},\mathbf{st_{3}}$, shown in Figure 2(a). All other states have the label $\mathbf{i}$, denoting intermediate states. The corresponding reward machine is shown in Figure 2(b). If the robot stacks the blocks in the order $\mathbf{st_{1}}\to\mathbf{st_{2}}\to\mathbf{st_{3}}$, it gets a reward of $1.0$. With a depth-$10$ prefix tree policy and $u_{\mathrm{max}}=3$, our algorithm recovers $\mathbf{2}$ consistent reward machines, which are the true reward machine up-to-renaming. The findings are summarized in Table 2. Results obtained from a finite set of demonstrations are reported in Appendix B.1.1.
@@
For our second task (stack-avoid), we introduce a “bad” state, shown in Figure 3(a). The true reward machine is shown in Figure 3(b). The robot’s task is to stack the blocks in the order $\mathbf{st_{1}}\to\mathbf{st_{2}}$ without going through $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}$}. If it does so, it reaches $u_{2}$ and gets a reward of $1$ forever. If during execution it passes through $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}$}, it will get a smaller (yet more immediate) reward of $0.2$, but it will get stuck at $u_{3}$ with 0 reward forever. We note that the product policy is uniformly random in both $u_{2}$ and $u_{3}$. This means that proposition traces such as $\mathbf{st_{1},i,st_{2}}$ and $\mathbf{st_{1},i,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{st_{bd}}},st_{2}}$ look identical from a policy perspective, as both reach nodes with uniformly random policies, while the first being more desirable than the second. By setting $u_{\mathrm{max}}=3$, a depth-$8$ policy was sufficient to recover the reward machine shown in Figure 3(c) up-to-renaming; that is, we find a smaller reward machine consistent for this task. Notably, setting $u_{\mathrm{max}}=4$ with the same depth policy yields more than a 1000 solutions, indicating that uniquely recovering the same sized reward machine -as the ground truth machine- requires longer depth policies. Results obtained from a finite set of demonstrations are reported in Appendix B.1.2.
@@
stack 10 60 9 73548 0.612
stack-avoid 8 60 9 24763 0.569
@@
5.3 2-Link Robotic Arm with Continuous State Space
Our setup for this experiment is a modified Reacher-v5 environment (Gymnasium (Towers et al., 2024); MuJoCo (Todorov et al., 2012)), where a planar robotic arm must reach targets randomly placed within the arena. The agent’s state is a 10‑D continuous vector consisting of the end‑effector’s position and velocity, the target’s coordinates, and their mutual distance. Actions are continuous torques in $1,1]^{2}$ applied at the elbow and shoulder joints.
Inspired by (Araki et al., 2021), we fix three colored targets — blue ($\mathbf{b}$), red ($\mathbf{r}$), and yellow ($\mathbf{y}$) — in distinct quadrants of the arena (see Figure 4). The goal is to reach them in the order $\mathbf{b}\to\mathbf{y}\to\mathbf{r}$. To train an expert policy, we discretize each torque dimension into five levels $\{-1.0,-0.5,0.0,0.5,1.0\}$ (yielding 25 total actions). We then employ Proximal Policy Optimization (PPO) (Schulman et al., 2017), as implemented in Stable‑Baselines3 (Raffin et al., 2021), to maximize a reward given by the negative Euclidean distance between the end‑effector and the active target.
@@
Table 3 reports, for various probability thresholds $\alpha$ (defined in Section 4.2), the size of the negative‐example set $\mathcal{E}^{-}$ and its False Positive Rate (FPR) [^1] FPR is simply the number of false positives divided by $|\mathcal{E}^{-}|$.. As expected from Section 4, solving our SAT problem with all the negative examples yields $\mathbf{0}$ solutions. Instead, we employ our weighted MAX-SAT variant and recover two viable solutions (Figures 4(b) and 4(c)) up-to-renaming, which represent the best achievable given our dataset. These are the same solutions that we recover if instead we supervised with the ground-truth reward machine to remove the false positives, emphasizing the robustness of this approach.